OpenStudy (anonymous):

Given the linear equation y=3/4x-2 , find the y-coordinates of the points (-8, ), (-4, ), and (4, ). Please show all of your work. Plot those points and graph the linear equation.

OpenStudy (anonymous):

simply put your x coordinates into the equation to get the y points

OpenStudy (anonymous):

sets are given in (x,y) for example the first set you have is (x,y)=(-8,y) for y=3/4x-2

OpenStudy (anonymous):

so for the first set x=-8 y=3/4(-8)-2=

OpenStudy (anonymous):

Given the linear equation , find the of the points so, the are (-8, -8 ), (-4, -5 ), and (4,1 ) check if these points lie in a same line